An athlete runs 105 miles south and 80miles west. How ,far is he from the starting point.
Answer by checkley77(12844)

105^2+80^2=C^2
11,025+6,400=C^2
17,425=C^2
C=SQRT17,425
C=132.004 MILES. ANS. FOR THE DISTANCE FROM THE STARTING POINT.